# Why do Machine Learning Notebooks Crash?

### v1

@@ 4.2 RQ2: To what extent are crashes in ML notebooks caused by ML bugs compared with general Python bugs?
Crashes in GitHub and Kaggle notebooks can stem from either code that uses ML libraries or general Python code. We analyze the distribution of ML/Python bugs to understand ML-specific crashes. Additionally, we examine the root causes of both ML and Python bugs, as shown in Figure 5, to gain deeper insights into their distinctions and similarities.
ML bugs and Python bugs
@@
ML bugs account for 64% of crashes overall in ML notebooks, with 59% on GitHub and 71% on Kaggle. The $\chi^{2}$ test reveals a significant difference in the distribution of ML and Python bugs between the two platforms ($p<0.001$). These findings suggest that a substantial proportion of crashes on both platforms stems from code that directly interacts with ML libraries, with the impact being particularly pronounced in Kaggle notebooks.
@@
Finding 2.1: A higher percentage of crashes in both GitHub and Kaggle notebooks are attributed to ML bugs (GitHub: 59%, Kaggle: 71%) compared to Python bugs. Kaggle exhibits an even more significant prevalence of ML-related crashes than GitHub.
@@
As Figure 5 illustrates, for ML-related crashes, the primary causes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%). Environment setup problems (12%) and implementation errors (10%) are also big factors. In contrast, general Python crashes are mainly driven by environment issues (24%) and implementation errors (24%), followed by notebook-specific issues (17%) and API misuse (13%).
Finding 2.2: The primary causes of ML-related crashes are API misuse (23%), data confusion (23%), and notebook-specific issues (21%), while general Python crashes are mostly caused by environment problems (24%), implementation errors (24%), and notebook-specific issues (17%).
@@
The root cause distributions between ML and Python bugs differ significantly ($p<0.001$). Notably, the higher percentage of API misuse in ML-related crashes (23%) compared to general Python crashes (13%) underscores the challenges developers face in correctly using ML libraries.
@@
Finding 2.3: API misuse is more prevalent in ML-related crashes (23%) than in general Python crashes (13%).
4.3 RQ3: In which stages of the ML pipeline are crashes in notebooks most likely to occur?
The ML pipeline outlines key stages of ML programs. We analyze the frequency of crashes occurring in these stages during ML development in notebooks and determine how often crashes in a certain stage involve ML bugs versus general Python bugs. The results are shown in Figure 6.
@@ 4.3 RQ3: In which stages of the ML pipeline are crashes in notebooks most likely to occur?
6(a) shows that most crashes occur during the data preparation (27% overall; GitHub: 29%, Kaggle: 26%) stage, model training (16% overall; GitHub: 13%, Kaggle: 19%), evaluation/prediction (15% overall; GitHub: 11%, Kaggle: 21%), and data visualization (11% overall; GitHub: 10%, Kaggle: 12%) stages.
Finding 3.1: The majority of crashes occurs during data preparation (27%), model training (16%), and evaluation/prediction (15%) stages in the ML pipeline.
Notably, 25% of crashes on GitHub are unrelated to any pipeline stage, which is approximately 3.5 times more than Kaggle (7%). We apply the $\chi^{2}$ test to evaluate this, revealing a significant difference in distributions ($p<0.001$).
@@
Finding 3.2: GitHub notebooks show 3.5 times more crashes than Kaggle not related to ML pipeline stages.
We then examine differences between GitHub and Kaggle in ML pipeline stages, excluding crashes unrelated to any stage. A $\chi^{2}$ test reveals a significant difference ($p<0.01$). To understand which stages contribute more to this difference, we calculate the adjusted residuals [34], comparing observed crashes in each stage with the expected number of crashes under the null hypothesis (shown in 6(b)). A larger absolute residual indicates a greater contribution. The results suggest that GitHub contains more crashes in the environment setup and data preparation, and Kaggle in the evaluation/prediction stage.
Finding 3.3: GitHub and Kaggle crashes are distributed differently across ML pipeline stages, especially in environment setup, data preparation, and evaluation/prediction.
@@
The breakdown of ML and Python bugs in each pipeline stage reveals that ML bugs dominate critical stages, with over 80% of crashes during model training and construction and over 70% during evaluation/prediction for both platforms.
Finding 3.4: ML-related crashes are dominant (over 70%) in model training, construction, and evaluation/prediction stages of the ML pipeline.
We apply Barnard’s exact test to assess significant differences of ML/Python bug distributions between GitHub and Kaggle for each ML pipeline stages. The results show significant differences in the environment setup ($p<0.05$) and data preparation ($p<0.01$) stages. In Kaggle, 70% of crashes during data preparation are ML-related, whereas in GitHub, ML and Python bugs occur at similar rates. In the environment setup stage, most Kaggle crashes are due to Python bugs, only 16% are ML bugs, while in GitHub, 45% of crashes involve ML bugs.
Finding 3.5: Kaggle has 1.4 times more ML-related crashes in the data preparation stage than GitHub, whereas GitHub shows almost 3 times more in the environment setup stage.
@@
We further analyze root cause distribution across ML pipeline stages (see Figure 7). Most stages display a range of root causes, including API misuse, notebook-specific issues, data confusion, environment issues, and implementation errors. Notably, in the environment setup stage, 84% of crashes stem specifically from environment-related issues.
@@ 5.1 Practical implications
As discussed in RQ1 (Section 4.1), notebook-specific issues are a major cause (19%) of crashes, contributing to 56% of NameError and 20% of KeyError exceptions. This highlights how notebook semantics can become a liability, increasing the risk of crashes. For instance, if a variable x is defined in one cell but a later cell that references x is executed first, a NameError occurs due to out-of-order execution since x has not been initialized yet. Similarly, if a developer incorrectly tries to add a column to a dataframe, causing the cell to crash, and then runs another cell that references this non-existent column without first correcting the initial error, a KeyError will occur due to previous cell error. The flexibility to execute cells in any order can easily disrupt cell dependencies, especially in complex ML notebooks. This underscores the unique challenges of bug detection and debugging in notebooks, which requires careful tracking of variable states and cell execution order at the current execution point. NBSAFETY [35] addresses this issue by providing a custom Jupyter kernel that flags potentially unsafe cell executions.
Takeaway 1: The flexibility of the notebook environment can cause instability, as out-of-order execution often leads to crashes like NameError.
@@
Another common root cause of crashes is data confusion (18%), indicating practical challenges in handling heterogeneous data types. It is a key contributor to a variety of exceptions such as ValueError, TypeError, KeyError, and IndexError, indicating that tackling data confusion could help resolve many of these exceptions. Data confusion also accounts for a high percentage of the ML-related crashes (23%), as discussed in RQ2 (Section 4.2). It suggests that improper data handling such as incorrect data manipulations for ML APIs, is critical in ML-specific code. Specifically, ML development often requires transformations of data into specific structures that can be easily mishandled, causing errors like shape mismatches, improper data transformations, or incorrect assumptions about data formats. This underscores the need for improved data validation.
Takeaway 2: Data handling poses a big challenge when using ML notebooks.
@@
When analyzing the results for RQ2 (Section 4.2), we find that crashes in ML notebooks are overall 1.8 times more frequently linked to ML bugs compared to Python bugs, reflecting key differences between developing ML and traditional software. Notably, a significantly higher proportion of ML-related crashes (23%) are directly attributed to API misuse, compared to only 13% for general Python crashes. The extensive reliance on ML libraries increases API misuse errors, particularly when developers are unfamiliar with libraries’ specific requirements. Although these libraries are crucial for effective data processing and modeling, they often have complex usage guidelines [36, 37]. Misinterpretations of parameter settings, or input-output expectations frequently result in crashes. To mitigate these issues, methods that verify API usage and enhance error-handling mechanisms could be beneficial.
Takeaway 3: Correct use of ML libraries is challenging in ML notebook development, highlighting the importance of verifying API usages and enhancing error-handling.
@@
RQ3 (Section 4.3) demonstrates that most ML notebook crashes occur during the data preparation stage (GitHub: 29%, Kaggle: 26%) of the ML pipeline. This highlights the primary challenge developers face: preparing or preprocessing datasets to ensure compatibility with model requirements. Issues often arise from understanding the data itself or aligning it with API expectations. Additionally, the limited support available for developers during this stage implies they may spend a disproportionate amount of time here. These insights suggest that improving tooling and resources for data preparation could significantly enhance developer productivity.
@@
Crashes are more common in the model training (16%) and evaluation/prediction (15%) stages than in the model construction (7%) stage. These errors typically arise at runtime, when models are trained or evaluated with actual data inputs, rather than during the construction phase when input shapes are often flexible. Libraries like TensorFlow, for instance, may not fully validate data compatibility until training begins, which can lead to errors surfacing only in these later stages. When issues occur within specific model layers, error messages can be ambiguous, making it difficult for developers to pinpoint and resolve problems. Consequently, developers may need to iteratively debug by re-training and re-evaluating multiple times, a tedious and time-intensive process.
Notebook semantics help preserve previous cell executions, so developers can avoid re-running all preceding cells (e.g., data preparation). However, this approach also increases the risk of crashes from, for example, out-of-order execution, where dependencies may no longer align. These challenges highlight the importance of enabling early bug detection without requiring code execution for training or evaluating ML models, along with mechanisms to track cell dependencies and maintain the correct execution order.
@@ 5.2 Comparison between GitHub and Kaggle
Comparing ML notebook crashes between GitHub and Kaggle reveals their key differences. GitHub has a significantly higher proportion of crashes outside typical ML pipeline stage (25% vs. 7% on Kaggle), suggesting a more diverse use of notebooks where crashes often occur outside standard ML workflows. Moreover, there is a notable difference in crash distributions across ML pipeline stages between the two platforms. For example, GitHub has more crashes during environment setup, likely due to its user-driven environments that can lead to inconsistencies in software versions or dependencies. GitHub also sees more crashes in data preparation, likely because it often uses “raw” datasets that require more complex preprocessing, whereas Kaggle provides standardized datasets with detailed descriptions. In contrast, Kaggle has more crashes during evaluation/prediction, possibly due to its submission process for competitions, which introduces additional model evaluation complexities.
@@
Kaggle shows a significantly higher prevalence of ML-related crashes than GitHub (71% vs. 59%). One exception is during the environment setup stage of the ML pipeline, where GitHub shows 3 times more ML-related crashes. This is likely because Kaggle’s pre-configured environment includes essential ML libraries, reducing setup-related errors. In contrast, GitHub, as a general-purpose code repository, preserves crashes from various environments. During the data preparation stage, Kaggle notebooks face 1.4 times more ML-related crashes, suggesting that Kaggle users focus more on ML-specific tasks, while GitHub users encounter a broader range of coding challenges. Furthermore, Kaggle provides a wealth of datasets alongside its notebooks, making it a valuable resource for reproducing experiments. However, existing studies [2, 38, 1, 6, 8] on notebooks have largely overlooked it as a data source.

### v2

@@ 4.2 RQ2: To what extent are crashes in ML notebooks caused by ML bugs and by which ML libraries?
Crashes in GitHub and Kaggle notebooks can stem from either code that uses ML libraries or general Python code. We analyze the distribution of ML/Python bugs to understand ML-specific crashes. Additionally, we examine the root causes of both ML and Python bugs, as shown in Figure 5, to gain deeper insights into their distinctions and similarities. Furthermore, we examine the frequency of crashes directly linked to the use of ML libraries (see Figure 6) to identify which libraries present the greatest challenges.
ML bugs vs. Python bugs
@@
ML bugs account for 67.7% of crashes overall in ML notebooks (GH: 62.8%, KG: 73.0%). The $\chi^{2}$ test reveals a significant difference in the distribution of ML and Python bugs between the two platforms ($p<0.01$). These findings suggest that a substantial proportion of crashes on both platforms stems from code that interacts with ML libraries, with the impact being particularly pronounced in Kaggle notebooks.
@@
Finding 2.1: On both platforms, we find significantly more ML bugs than Python bugs. Kaggle exhibits a more significant prevalence of ML-related crashes than GitHub.
@@
As Figure 5 illustrates, the primary causes of ML-related crashes are API misuse (25%), notebook-specific issues (20%), and data confusion (20%). Environment setup problems (12%) and implementation errors (11%) are also big factors. Furthermore, ML-specific bugs include unique root causes such as ML model mismatch (3%), insufficient resources (3%), and library issues (1%). In contrast, general Python crashes are mainly driven by implementation errors (29%) and environment issues (28%), followed by notebook-specific issues (18%) and API misuse (13%). The differences in root causes between ML and general Python bugs stem from the inherent complexity of working with ML models and data, as well as the common use of notebook semantics when performing ML tasks. While the saved states in notebooks improve usability, they can also increase the likelihood of crashes if not properly managed.
Finding 2.2: The primary causes of ML-related crashes are API misuse (25%), notebook-specific issues (20%), and data confusion (20%), while general Python crashes are mostly caused by implementation errors (29%) and environment problems (28%).
@@
The root cause distributions between ML and Python bugs differ significantly ($p<0.001$). Notably, the higher percentage of API misuse in ML-related crashes (25%) compared to general Python crashes (12%) underscores the challenges developers face in correctly using ML libraries.
@@
Finding 2.3: API misuse is more prevalent in ML-related crashes (25%) than in general Python crashes (12%).
Figure 6: The challenge factors ($CF$) of the top 10 ML libraries, indicating the ratio between a library’s involvement in crashes and its usage. See Table II for library descriptions.
ML libraries
Figure 6 ranks the top 10 most commonly used ML libraries by their challenge factor ($CF$), which measures how likely using a ML library is to cause crashes compared to how often it is used. The Challenge factor is computed as $CF=Lib_{\text{crashes}}/Lib_{\text{usage}}$, where 1) $Lib_{\text{crashes}}=N_{\text{Lib\_crash}}/N_{\text{crashes}}$ is the proportion of crashes caused by using an ML library. 2) $Lib_{\text{usage}}=N_{\text{Lib\_use}}/N_{\text{notebooks}}$ represents how frequently the library appears in ML notebooks, counting each notebook only once per library. A higher $CF$ indicates a library is more crash-prone relative to its usage. In this ranking, we merge tensorflow and keras because keras has been tightly integrated into tensorflow since TensorFlow 2.0. Combining them provides a more accurate representation of their impact compared to other ML libraries. ML libraries tensorflow/keras, and torch exhibit the highest challenge factors (over 5), followed by sklearn (2.8), pandas (2.5), and torchvision (1.3).
Finding 2.4: tensorflow/keras and torch are the top ML libraries most challenging to use correctly.
4.3 RQ3: In which stages of the ML pipeline are crashes in notebooks most frequently to occur?
The ML pipeline outlines the key stages of ML programs. We analyze the frequency of crashes occurring in these stages during ML development in notebooks and determine how often crashes in a certain stage involve ML bugs. The results are shown in Figure 7.
@@ 4.3 RQ3: In which stages of the ML pipeline are crashes in notebooks most frequently to occur?
7(a) shows that most crashes occur during the data preparation (33.0% overall; GH: 37.7%, KG: 27.8%) stage, model training (19.4%; GH: 17.4%, KG: 21.6%), evaluation/prediction (18.4%; GH: 14.6%, KG: 22.5%), and data visualization (13.5%; GH: 14.1%, KG: 12.9%) stages. Crashes during data preparation are likely due to the complexities of cleaning and processing datasets. Model training and evaluation crashes may result from incorrect settings, misaligned inputs, or resource limitations. Data visualization errors are likely caused by mismatched data types or improper library usage.
Finding 3.1: The majority of crashes occurs during data preparation (33.0%), model training (19.4%), and evaluation/prediction (18.4%) stages in the ML pipeline.
@@
We then examine the differences between GitHub and Kaggle in the ML pipeline stages. A $\chi^{2}$ test reveals a significant difference ($p<0.01$). To understand which stages contribute more to this difference, we calculate the adjusted residuals [39], comparing observed crashes in each stage with the expected number of crashes under the null hypothesis (shown in 7(b)). A larger absolute residual indicates a greater contribution. The results suggest that GitHub contains more crashes in the environment setup and data preparation, likely due to the variability in user environments and dependencies. Kaggle, with pre-configured environment, faces more crashes during evaluation/prediction, possibly due to model or input setting issues, or resource limitations.
Finding 3.2: GitHub and Kaggle crashes are distributed differently across ML pipeline stages. GitHub has more during environment setup and data preparation, whereas Kaggle sees more during the evaluation/prediction stage.
@@
The breakdown of ML and Python bugs in each pipeline stage reveals that ML bugs dominate critical stages, with over 80% of crashes during model construction and training, and over 70% during evaluation/prediction for both platforms.
Finding 3.3: ML-related crashes are dominant (over 70%) in model construction, training, and evaluation/prediction stages of the ML pipeline.
We apply Barnard’s exact test to assess significant differences of ML/Python bug distributions between GitHub and Kaggle for each ML pipeline stage. The results show significant differences in the data preparation ($p<0.01$) stage. In Kaggle, 70% of crashes during data preparation are ML-related, whereas in GitHub, ML and Python bugs occur at similar rates.
Finding 3.4: Kaggle has 1.4 times more ML-related crashes in the data preparation stage than GitHub.
@@
We further analyze root cause distribution across ML pipeline stages (see Figure 8). Most stages display a range of root causes, including API misuse, notebook-specific issues, implementation errors, environment issues, and data confusion. Notably, in the environment setup stage, 84% of crashes stem specifically from environment-related issues.
@@ 5.1 Practical implications
As discussed in RQ1 (Section 4.1), notebook-specific issues contribute to a significant portion (19.4%) of crashes, particularly those related to name resolution (e.g., Variable Not Found, Name Error, Module Not Found), Model Initialization Error, and key errors. These results highlight how the flexible execution model of notebooks can introduce instability, increasing the likelihood of crashes.
Unlike conventional software engineering workflows, where execution order is typically linear and well-defined, notebooks enable an interactive, fragmented development process. This flexibility, while beneficial for experimentation, also disrupts state management, leading to errors caused by out-of-order execution and implicit dependencies. For instance, if a variable x is defined in one cell but a later cell that references x is executed first, a Variable Not Found occurs due to out-of-order execution since x has not been initialized yet. These issues underscore the unique challenges of debugging in notebooks, where errors may not be immediately apparent and can propagate unpredictably.
This finding underscores notebook development challenges, particularly in managing hidden state and execution order dependencies. Tools like NBSAFETY [40] attempt to mitigate these risks by providing a custom Jupyter kernel that detects unsafe execution sequences. However, our results suggest that the problem remains prevalent, reinforcing the need for improved debugging support tailored to the dynamic execution model of notebooks.
Takeaway 1: The flexibility of notebooks disrupts state management, leading to unusual execution-order errors, which highlight the need for better debugging tools to support ML developers.
@@
Another common root cause of crashes is data confusion (16.4%), indicating practical challenges in handling heterogeneous data types. There are many crash types, such as Key Error, Data Value Violation, Unsupported Broadcast, and Index Error, which point to the same root cause of data confusion. This potentially complicates fault localization process. Data confusion also accounts for a high percentage of ML-related crashes (20%) (see Section 4.2), emphasizing the impact of improper data handling, such as incorrect transformations for ML APIs. Specifically, ML development often requires precise data formatting, and errors like shape mismatches or faulty transformations can be hard to allocate. This underscores the need for better data validation tools to prevent such issues.
Takeaway 2: Data confusion in ML notebooks leads to diverse crash types, making fault localization challenging.
@@
When analyzing the results for RQ2 (Section 4.2), we find that crashes in ML notebooks are overall 2.1 times more frequently linked to ML bugs compared to Python bugs, reflecting key differences between developing ML and traditional software. Notably, a significantly higher proportion of ML-related crashes (25%) are attributed to API misuse, compared to only 13% for general Python crashes.
This issue stems from the heavy reliance on ML libraries, with complex usage guidelines [41, 42] that often lead to misinterpretations of parameters and input-output expectations, causing crashes. Our analysis reveals that certain DL libraries, particularly tensorflow/keras and torch, pose greater challenges. In tensorflow/keras, tensor shape mismatches are a major cause of crashes, often due to implicit broadcasting, dynamic input shapes, or insufficient safeguards for detecting misaligned dimensions before execution. In contrast, torch is more prone to runtime errors, such as device compatibility issues (e.g., loading CUDA models on a CPU) and incorrect loss function configurations. Some of these errors may arise from unclear documentation, while others may stem from the flexibility of these libraries, where errors only surface at runtime, making static checks difficult. More informative error messages and warnings could significantly aid debugging.
These findings suggest that many crashes may stem not just from user mistakes but also from library design choices, including complex APIs, ambiguous error messages, and limited safeguards. This highlights the need for library developers to improve API usability, enhance documentation clarity, and implement better error-handling mechanisms. Additionally, identifying error-prone libraries can help practitioners prioritize debugging efforts, develop better tools, and choose libraries that align with their stability needs.
Takeaway 3: ML notebooks are prone to crashes due to API misuse, especially when using deep learning libraries tensorflow/keras and torch.
@@
RQ3 (Section 4.3) demonstrates that ML notebook crashes occur mostly during the data preparation stage (33.0%) of the ML pipeline. This highlights the primary challenge developers face: preparing or preprocessing datasets to ensure compatibility with model requirements. Issues often arise from understanding the data itself or aligning it with API expectations. Additionally, the limited support available for developers during this stage implies that they may spend a disproportionate amount of time here. These insights suggest that improving tooling and resources for data preparation could significantly enhance developer productivity.
@@
Crashes are more common in the model training (19.4%) and evaluation/prediction (18.4%) stages than in the model construction (8.3%) stage. These errors typically arise at runtime, when models are trained or evaluated with actual data inputs, rather than during the construction phase, when input shapes are often flexible. Libraries like tensorflow/keras, for instance, may not fully validate data compatibility until training begins, which can lead to errors surfacing only in these later stages. When issues occur within specific model layers, error messages can be ambiguous, making it difficult for developers to pinpoint and resolve problems. Consequently, developers may need to iteratively debug by re-training and re-evaluating multiple times, a tedious and time-intensive process.
Developers can avoid re-running all preceding cells as notebook semantics preserve previous cell executions, speeding up the iterative debugging process. However, this also increases the risk of crashes from, for example, out-of-order execution, where dependencies may no longer align. These challenges highlight the importance of enabling early bug detection without requiring code execution for training or evaluating ML models, along with mechanisms to track cell dependencies and maintain the correct execution order.
@@ 5.2 Comparison between GitHub and Kaggle
Comparing ML notebook crashes between GitHub and Kaggle reveals their key differences. In the cases excluded from the initial sample, GitHub has a significantly higher proportion of crashes outside typical ML pipeline stages (25% vs. 7% on Kaggle), for example, tutorial notebooks, educational materials, or physics simulations. This suggests a more diverse use of notebooks where crashes often occur outside standard ML workflows. Moreover, there is a notable difference in crash distributions across ML pipeline stages between the two platforms. For example, GitHub sees more crashes in data preparation, likely because it often uses “raw” datasets that require more complex preprocessing, whereas Kaggle provides standardized datasets with detailed descriptions. In contrast, Kaggle has more crashes during evaluation/prediction, possibly due to its submission process for competitions, which introduces additional model evaluation complexities.
@@
Kaggle shows a significantly higher prevalence of ML-related crashes than GitHub (73.0% vs. 62.8%). One exception is during the environment setup stage of the ML pipeline, where GitHub shows 2.6 times more ML-related crashes. This is likely because Kaggle’s pre-configured environment includes essential ML libraries, reducing setup-related errors. In contrast, GitHub, as a general-purpose code repository, preserves crashes from various environments. During the data preparation stage, Kaggle notebooks face 1.4 times more ML-related crashes, suggesting that Kaggle users focus more on ML-specific tasks, while GitHub users encounter a broader range of coding challenges. Furthermore, Kaggle provides a wealth of datasets alongside its notebooks, making it a valuable resource for reproducing experiments. However, existing studies [2, 43, 1, 6, 10] on notebooks have largely overlooked it as a data source.